OpenStudy (jamesreed30):

Find the approximate value of the circumference of a circle with the given radius use PI= 3.14 round your answer to the nearest thousandth. 3.21 ft

OpenStudy (jiteshmeghwal9):

circumference of a circle = 2 \(\pi\) r r=3.21 ft \(\pi=3.14\)

OpenStudy (ovidia):

formula for circumference: P=2πr P=2*3.14*4.6 P=28.888 in= 28.89 in

OpenStudy (jamesreed30):

So 3.14 is the answer?

OpenStudy (jamesreed30):

Or 28.89 is

OpenStudy (ovidia):

what do you think is the answer?

OpenStudy (jamesreed30):

28.89?

OpenStudy (ovidia):

that's what id go with tell me if it was right

OpenStudy (jamesreed30):

Can you do one more for me please? ITS the same question but the numbers are now 6.53

OpenStudy (ovidia):

yes

OpenStudy (ovidia):

@jamesreed30

OpenStudy (jamesreed30):

Okay so if it's 6.53 cm what is it?

OpenStudy (ovidia):

32.03

OpenStudy (ovidia):

what did you get?

OpenStudy (jamesreed30):

It said they weren't right

OpenStudy (ovidia):

both? dang

OpenStudy (jamesreed30):

Yeah it's alright though lol